OpenStudy (anonymous):

What is the solution to the equation 4^(-2x)=3? x= -.631 x=-.396 x=.396 x=.631

OpenStudy (nurali):

\[\frac{ 1 }{ 4^{2x} }=3\] \[\frac{ 1 }{ 3 }=4^{2x}\] \[0.33=4^{2x}\] Log both sides \[\log(0.33)=\log(4^{2x})\] \[-0.48=2xlog(4)\] \[-0.48=2x(0.60)\] \[-0.48=1.20x\] \[x=-0.4\]

OpenStudy (anonymous):

that's not one of the answers?